I am also having a problem with this.
The diameter of the Milky Way disc is approximately 9x10^20 meters. How lon does it take light, travel at 10^16 m/year to travel across the diameter of the Milky Way?
: I am also having a problem with this.
The diameter of the Milky Way disc is approximately 9x10^20 meters. How lon does it take light, travel at 10^16 m/year to travel across the diameter of the Milky Way?

Answer by edjones(2391)

9*10^20/10^16=9*10^4=90,000 yrs